SCIENTIFIC PROGRAMME

Thursday, 11

10:30 – 11:30 R. Osserman (invited lecture) A new variant of the Schwarz–Pick–Ahlfors Lemma

11:30 – 12:00 E. García–Río (session B_{1}) On a problem of Osserman concerning the Jacobi operator

J. C. González-Dávilla (session B_{2})

Geodesic Spheres, Tubes and Flow Space Forms

12:00 – 12:30 L. Ugarte (session B_{1}) The holomorphic fiber bundle structure of some compact complex nilmanifolds

A. M. Azevedo Breda (session B_{2})

Monohedral planar f-tilings

14:30 – 15:00 I. da Silva (session B_{1}) An Euclidean Characterization of the Regions of an Arrangement of Pseudospheres

J. Perez (session B_{2})

Strong rigidity and periodic minimal surfaces

15:00 – 15:30 A. Guedes de Oliveira (session B_{1}) An interpretation of the monodromy group of a wiring diagram

F. Martin (session B_{2})

Huge Singly Periodic Minimal Surfaces in R^{3} and Symmetries

15:30 – 16:00 Poster Presentation Pedro Albuquerque, Ester Sanabria Codesal and Marta Macho-Stadler

17:00 – 17:30 J.A. Armario (session B_{1}) On the computability of the p-local homology of twisted Cartesian products of Eilenberg-MacLane spaces

S. Carter (session B_{2})

On k-minimal submanifolds

17:30 – 18:00 J. Rodriguez (session B_{1}) On cellular approximations using Moore Spaces

I. Castro (session B_{2})

Minimal Lagrangian submanifolds in complex Euclidean space Friday, 12

09:30 – 10:00 H. R. Farran (session B_{1}) Smooth exact fillings

A. Ferrandez (session B_{2})

Solutions of the Betchov-Da Rios soliton equation: a Lorentzian approach

10:00 – 10:30 B. Wegner (session B_{1}) Variation of the Total Normal Twist of Closed Curves in Euclidean Spaces

J. M. Nunes da Costa (session B_{2})

Bi-Hamiltonian manifolds – an application to the relativistic Toda lattice and a generalization to the Jacobi manifolds

10:30 – 11:00 F. Craveiro de Carvalho (session B_{1}) The parallel group of a plane curve

R. Ibañez (session B_{2})

Some Generalizations of Poisson and Jacobi Structures

11:30 – 12:30 D. Salamon (invited lecture) Stable maps: Gromov compactness and the Arnold conjecture

14:30 – 15:30 W. Ballmann (invited lecture) Spaces of nonpositive curvature

17:00 – 17:30 A. Tarrio-Tobar (session B_{1}) Some results about Integral Geometry in the n-dimensional quaternionic space

L. J. Alias (session B_{2})

Mean isotropic stability of zero mean curvature surfaces in spacetimes

17:30 – 18:00 I. Sánchez-Rodríguez (session B_{1}) Explicit Formulas for the 3-Jet Lift of a Matrix Group. Apllications to Conformal Geometry

M. C. Romero-Fuster (session B_{2})

Some Global Properties of Surfaces in 4-Space Saturday,12

10:30 – 11:00 R. M. B. Chaves (session B_{1}) On two parameter envelopes of N-spheres in a real space form

M. E. Galvão (session B_{2})

Mean Curvature One Surfaces in Hyperbolic Space

11:30 – 12:30 A. Verjovsky (invited lecture) Examples of higher diimensional Kleinian groups